OpenStudy (anonymous):

Hi! I'm working on my homework. I need a little help with this question. Does someone mind helping? You have 1200 feet of fencing and you want to make two fenced in enclosures by splitting one enclosure in half. What are the largest dimensions of this enclosure that you could build? 125 ft x 200 ft 100 ft x 105 ft 300 ft x 200 ft 105 ft x 120 ft

OpenStudy (shirley27):

First you have to make an equation for this question, do you have an idea?

OpenStudy (anonymous):

Yes. I know that 3x+2y=1200

OpenStudy (shirley27):

w = width and l = length 3w + 2l = 1200 w*l = area If you isolate the first equation for l, you get: l = 600 - 3/2w Then you could plug this into the second equation to get a parabolic formula. By finding the maximum value through either graphing or completing the square, you could find the maximum area.

OpenStudy (anonymous):

okay... im a little lost... do i need my graphing calculator? i know ive done this without one before...

OpenStudy (misssunshinexxoxo):

Are you familiar with area?

OpenStudy (shirley27):

I think i've complicated things a little

OpenStudy (anonymous):

yes. i kinda know how to do this, but all my ways have failed

OpenStudy (misssunshinexxoxo):

Be confident I'll walk you through this

OpenStudy (anonymous):

okay

OpenStudy (misssunshinexxoxo):

do you have a calculator?

OpenStudy (misssunshinexxoxo):

tell me what is 300 x 300?

OpenStudy (anonymous):

90,000

OpenStudy (misssunshinexxoxo):

ok we have 4 sides; the answer is C